OpenStudy (anonymous):

f(x)= 3x – 7 and g(x) = –2x – 6. Find (ƒ ∘ g)(4).

OpenStudy (mertsj):

f composite g means replace each x in the function f with the function g. So in your example, take the function f and replace each x with -2x-6

OpenStudy (anonymous):

i still don't get it..

OpenStudy (mertsj):

What is the function f?

OpenStudy (anonymous):

i don't know

OpenStudy (mathstudent55):

(ƒ ∘ g)(4) = f(g(4)) This means that the composition of function f and function g means to evaluate function g at 4, g(4), then insert that value into function f, f(g(4)).

OpenStudy (mertsj):

Read the problem if you want to know what the function f is.

OpenStudy (mathstudent55):

g(x) = -2x - 6 g(4) = -2(4) - 6 Find what g(4) is equal to. Then take that value and insert it in for x of f(x) = 3x - 7

OpenStudy (anonymous):

insert it where

OpenStudy (mertsj):